OpenStudy (konradzuse):

The displacement from equilibrium of a mass oscillating on the end of a spring suspended from a ceiling is given below, where y is the displacement in feet and t is the time in seconds. Use a graphing utility to graph a displacement function on the interval [0,10]. Find the value of t past which the displacement is less than 3 inches from equilibrium. (Round your answer to two decimal places.) y = 1.7e^(−0.22t) * cos(4.9t) t = ??? seconds

OpenStudy (konradzuse):

Anyone? :(

OpenStudy (anonymous):

substitute y by 3 and solve for t. 3 = 1.7e^(−0.22t) * cos(4.9t)

OpenStudy (konradzuse):

Hmm....

OpenStudy (asnaseer):

the maximum displacements will be achieved when the cos() term is equal to 1. this happens when:\[4.9t=n\pi\]therefore, the maximum displacements occur when:\[t=\frac{n\pi}{4.9}\]

OpenStudy (konradzuse):

so in this case n = 3?

OpenStudy (asnaseer):

substitute that into your equation and set y=3 to get:\[3=1.73^{-0.22\times\frac{n\pi}{4.9}}\]

OpenStudy (asnaseer):

solve this for n, which will then give you a value for t

OpenStudy (asnaseer):

sorry - replace 1.73 by 1.7e above

OpenStudy (konradzuse):

I keep feeling like we need a log to bring down the n?

OpenStudy (asnaseer):

\[3=1.7e^{-0.22\times\frac{n\pi}{4.9}}\]

OpenStudy (asnaseer):

yes - you will need to use logs

OpenStudy (konradzuse):

sweet okay.

OpenStudy (asnaseer):

also, you can use wolframalpha to graph this. see here: http://www.wolframalpha.com/input/?i=y+%3D+1.7e^%28%E2%88%920.22t%29+*+cos%284.9t%29+for+t%3D0+to+10

OpenStudy (konradzuse):

ln 3 = ln -0.22 + ln(n*pi) +ln (1.7e) ?

OpenStudy (konradzuse):

not sure if my log rules are correct...

OpenStudy (asnaseer):

first divide both sides by 1.7 to get:\[\frac{3}{1.7}=e^{-0.22\times\frac{n\pi}{4.9}}\]then take log to base e to get:\[\log(\frac{3}{1.7})=-0.22\times\frac{n\pi}{4.9}\]

OpenStudy (konradzuse):

ah okay. :)

OpenStudy (konradzuse):

0.568 is what log(3/1.7) is so should I keep the orignal, or this?

OpenStudy (konradzuse):

then we would need to multiply by 4.9 and divide by 0.22 pi

OpenStudy (asnaseer):

I always leave the actual calculations to the end to avoid any rounding errors.

OpenStudy (konradzuse):

okies.

OpenStudy (konradzuse):

so then we should have\[\frac{ 4.9 \log(3/1.7)}{0.22\pi}\]

OpenStudy (konradzuse):

= n

OpenStudy (asnaseer):

you missed the negative sign.

OpenStudy (konradzuse):

oh didn't even see that...

OpenStudy (konradzuse):

which = -4.03

OpenStudy (konradzuse):

so would it be 4.03 seconds?

OpenStudy (asnaseer):

yes - if you look at its graph you will see that at t=0, its displacement from the equilibrium is about 1.6. so this means t has to be negative for the displacement to be bigger than this. are you sure the question is stated correctly?

OpenStudy (konradzuse):

yup

OpenStudy (asnaseer):

hang on - we missed something!

OpenStudy (asnaseer):

y is said to be in feet. and we are supposed to find displacement of 3 /inches/!

OpenStudy (asnaseer):

so we should have set y = 3/12 = 1/4 = 0.25

OpenStudy (asnaseer):

so just replace the 3 in your final equation by 0.25

OpenStudy (konradzuse):

(4.9log(0.25/1.7))/(0.22pi) ?

OpenStudy (asnaseer):

yes

OpenStudy (konradzuse):

13.59

OpenStudy (asnaseer):

again - don't forget the minus sign!

OpenStudy (konradzuse):

yeah it was negative 13.59 :)

OpenStudy (asnaseer):

no the answer works out to be plus 13.59

OpenStudy (konradzuse):

hmm that's apparently still not right. Yeah it was negative 13.59 without the negative sign, so it's positive.

OpenStudy (asnaseer):

\[-\frac{4.9 \log(0.25/1.7)}{0.22\pi}\approx13.59\]

OpenStudy (asnaseer):

and remember this is the value for n

OpenStudy (asnaseer):

\[t=\frac{n\pi}{4.9}\]

OpenStudy (konradzuse):

t = npi/4.9

OpenStudy (konradzuse):

so t = 13.59pi/4.9

OpenStudy (asnaseer):

yes

OpenStudy (konradzuse):

8.71.

OpenStudy (asnaseer):

that looks right

OpenStudy (konradzuse):

still not right hmm.... FACEPALM.

OpenStudy (konradzuse):

I don't like this q :(

OpenStudy (asnaseer):

let me re-check everything above...

OpenStudy (konradzuse):

thanks :).

OpenStudy (konradzuse):

Is it because we forgot that cos?

OpenStudy (konradzuse):

But I guess that was the npi.

OpenStudy (asnaseer):

ok, I think when we found n, we needed to round it to the next WHOLE number as n is supposed to be an integer. so we should have used n=14, which would then give us:\[t=\frac{14\pi}{4.9}\approx 8.98s\]

OpenStudy (konradzuse):

Nope that doesn't work either :(

OpenStudy (asnaseer):

:/

OpenStudy (konradzuse):

!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (asnaseer):

the only other option I can think of is that we should have rounded n down to 13 instead which would give:\[t=\frac{13\pi}{4.9}\approx 8.33 s\]

OpenStudy (konradzuse):

according to this practice one this is the answer, maybe we will try to see if we can get the same 8.40? I'm not sure if it should be higher or lower because 1.66 is lower than 1.7..

OpenStudy (asnaseer):

maybe my assumption that the cos() term must equal 1 was flawed?

OpenStudy (konradzuse):

Idk I asked turning to come help :D. Maybe he would know... :P

OpenStudy (asnaseer):

hang on - from that image it looks like you can use the Newton-Raphson method

OpenStudy (asnaseer):

so just use that and iterate to the required level of accuracy

OpenStudy (konradzuse):

Never heard of that method..... I thought they were talking about one of newton's laws of physics :P

OpenStudy (asnaseer):

:)

OpenStudy (asnaseer):

I guess you are also allowed to approximate the answer by using your graphing utility of choice. i.e. just visually check when the displacement becomes less that 3" from the equilibrium by inspecting the graph

OpenStudy (konradzuse):

Sadly no graphing calculator, only wolfman! :)

OpenStudy (asnaseer):

try this: https://www.desmos.com/calculator

OpenStudy (turingtest):

shall I post what I think is the solution through wolfram?

OpenStudy (asnaseer):

TuringTest - I believe the solution is 8.40

OpenStudy (konradzuse):

8.40 is the solution to the practice :P.

OpenStudy (turingtest):

oh I missed that the solution was known, sorry

OpenStudy (konradzuse):

I onyl have 1 more chance to get it right so lets hope we can do this, if not an hour wasted ftl :(

OpenStudy (konradzuse):

Na that's the solution to the practice question using 1.66 instead of 1.7... Idk what the solution to this ugly one is :(

OpenStudy (konradzuse):

unless that's what we are talking ABOOOT :P.

OpenStudy (turingtest):

http://www.wolframalpha.com/input/?i=solve+0.25+%3D+1.7e%5E%28%E2%88%920.22t%29+*+cos%284.9t%29+for+0+to+10 this indicates to me that it is about 7.8

OpenStudy (asnaseer):

the effort here is not wasted @KonradZuse - it is more important to understand HOW to solve a problem than it is to get any particular answer correct.

OpenStudy (turingtest):

but that's just me playing with a computer, exactly^

OpenStudy (konradzuse):

Yeah I know :P. This has been an interesting learning experience :P.

OpenStudy (konradzuse):

http://www.wolframalpha.com/input/?i=solve+0.25+%3D+1.66e%5E%28%E2%88%920.22t%29+*+cos%284.9t%29+for+0+to+10 This produces the same thing abouts.

OpenStudy (asnaseer):

My gut feeling is to use the Newton-Raphson method to solve this numerically.

OpenStudy (konradzuse):

lets doo it then :P. LEts try the practice one first :).

OpenStudy (asnaseer):

but as you haven't learnt that method yet I guess you will have to stay with the graphical method

OpenStudy (asnaseer):

oh - so you want to learn that method?

OpenStudy (konradzuse):

is it calc 1 stuff? I basically forgot everything in calc 1 :P. Calc 2 is rough :P.

OpenStudy (asnaseer):

have you learnt about calculus yet?

OpenStudy (konradzuse):

All this stuff seems to be a revamp of calc 1....

OpenStudy (konradzuse):

I took calculus 1 online about 3 years ago.

OpenStudy (asnaseer):

the methos itself is described here: http://en.wikipedia.org/wiki/Newton%27s_method

OpenStudy (asnaseer):

basically we use the iterative approach:\[x_{n+1}=x_n-\frac{f(x_n)}{f'(x_n)}\]where f'(x) is the differential of f(x)

OpenStudy (asnaseer):

in your case we are given:\[0.25=1.7e^{−0.22t} *\cos(4.9t)\]so this needs to be re-arranged to get:

OpenStudy (asnaseer):

\[f(t)=0.25-1.7e^{−0.22t} *\cos(4.9t)\]and we are trying to find a value for t that gives a zero of f(t)

OpenStudy (asnaseer):

making sense so far?

OpenStudy (konradzuse):

mhm

OpenStudy (asnaseer):

so you need to first find the derivative of f(t). do you know how to do this?